# Intermittent localization and fast spatial learning by non-Markov random walks with decaying memory

### v1

@@ III Previous results for the perfect memory case
Inserting these expressions into Eq. (II) gives
@@
This nonlinear equation can be exactly solved for the Fourier transform of $P_{n}^{(st)}$, where the occupation probability of the origin $P_{0}^{(st)}$ is considered as a constant that is finally obtained self-consistently. The analytical expression for the occupation probability at the origin was obtained in [20] and reads,
@@
A quick analysis of Eq. (9) shows that $P_{0}^{(st)}>0$ for all $q>0$ and $\gamma>0$. Consequently, the localization length $\xi=(\ln a)^{-1}$ is always finite and the walker localizes for any non-zero parameter values $q$ and $\gamma$, a fact that was checked in numerical simulations of the model [20]. In addition, the agreement of Eqs. (9)-(10) with the numerical simulations was very good (as also illustrated by Fig. 1b below), suggesting that the de-correlation approximation above might be exact. With the solution (9)-(10) at hand, the participation ratio defined by
$PR=\sum_{n=-\infty}^{\infty}P_{n}^{2}$ (13)
is given by the expression
@@
Hence PRPR is finite, independent of the system size.
Localization is this context can be interpreted as a phenomenon of spatial learning, where the walker “learns” to exploit the resource patch at the origin instead of diffusing far away from it. As shown below in Fig. 1b, localization is stronger ($P_{0}^{(st)}$ larger) if the resource patch is more attractive ($\gamma$ large) or memory more frequently used ($q$ large). Interestingly, on $3d$ lattices, the same de-correlation approximation predicts that a localization/delocalization phase transition takes place at $q=q_{c}$ with $q_{c}$ a non-zero critical value, in agreement with simulations [21]. In addition, the localization length $\xi$ diverges near $q_{c}$ with the same critical exponents as in the self-consistent theory of Anderson localization [20, 21]. If the nearest-neighbor random walk is replaced by a Lévy flight with sufficiently small index, a localization/delocalization transition takes place on the $1d$ lattice, too [20]. In the problem studied here, $q_{c}=0$ and there is no de-localized phase.
@@
We now address the case of a general memory kernel $\pi_{t,t^{\prime}}$. If we apply the same approximations (5)-(III) to Eq. (II), one obtains Eq. (8) again. This is simply due to the fact that the normalization condition $\sum_{t^{\prime}=0}^{t}\pi_{t,t^{\prime}}=1$ is valid for all kernels. Hence, under the same assumptions, the results (9)-(14) hold for any kernel, not just the perfect memory case. This, of course, cannot be correct if memory is very short-range, with $\pi_{t,t^{\prime}}$ very peaked near the present time $t$, i.e., $F(\tau)$ tends rapidly to 0 in Eq. (2). In that case, one expects the walker to behave roughly as a standard random walk with a modified diffusion coefficient, as shown in [34] in the absence of impurity. No stationary localized solution should be expected in principle.
To built intuition on the behavior of the system in the different cases, we have run Monte Carlo simulations that generate stochastic trajectories obeying the rules of the model. Appendix A gives details on the generation of the random times $t^{\prime}$. We have considered the following choices for the function $F(\tau=t-t^{\prime})$.
Figure 1: (a) A localized trajectory $X_{t}$ vs. t with power-law memory decay, for $\beta=1$, obtained from a numerical simulation with $q=0.1$ and $\gamma=0.7$. (b) Occupation probability of the origin, $P_{0}$, as a function of $q$ for various values of $\gamma$. Solid (dark blue) lines are given by Eq. (9) and symbols by Monte Carlo simulations ($t=10^{6}$) corresponding to different memory decay exponents $\beta\in[0,1]$.
@@ IV.1 Slowly decaying memory
In the first example, the memory slowly decays as a power-law, and not faster than $1/\tau$. This type of law was observed in experiments on human subjects (the Ebbinghaus’ forgetting curve [42]) or inferred in the wild on large herbivores [32], for instance. We denote,
@@
where $\tau=0,1,$... In this range of exponent values, the mean forgetting time defined as [34]
$\langle\tau\rangle=\lim_{t\to\infty}\frac{\sum_{t^{\prime}=0}^{t}(t-t^{\prime})F(t-t^{\prime})}{\sum_{t^{\prime}=0}^{t}F(t-t^{\prime})}$ (16)
diverges. In addition, the effective number of visits that are remembered by the walker up to time $t$ [Eq. (3)] keeps increasing unbounded, albeit more slowly than $t$. Hence memory is quite long-ranged for these kernels and the perfect memory (non-decaying) is recovered with $\beta=0$.
We display in Fig. 1a a typical trajectory $X_{t}$ vs. $t$ with $\beta=1$, showing a clear stationary localized behavior at late times, where the particle is never far away from the origin. Figure 1b shows the occupation probability $P_{0}$ computed at $t=10^{6}$ as a function of $q$ for various values of $\gamma$. In each case, the curves corresponding to different values of $\beta$ in the interval $[0,1]$ practically fall on top of each other and are in very good agreement with the theoretical prediction (9) of the de-correlation approximation. Hence, the range of applicability of this approximation is not limited to the perfect memory studied in [20] and extends to kernels with sufficiently slow forgetting as well.
@@ IV.2 Exponentially decaying memory
Two typical trajectories are displayed in Fig. 2a-b and a new type of behavior can be observed: the walker is not completely localized nor diffusive but stochastically alternates between these two types of motion. Fig. 2b, for instance, shows a very long diffusive phase at late times. We have identified the two phases using a simple algorithm described in Appendix B and have analyzed them separately. Notably, the occupation probability $P_{0}$ at the origin computed from averaging only over the localized parts of many trajectories, shown in Fig. 2c, is practically undistinguishable from $P_{0}^{(st)}$ in Eq. (9). In comparison, the occupation probability $P_{0}$ in the de-localized phase is very low. (Note that if $q$ is too small, the identification of the two phases becomes difficult and cannot be done in a reliable way; these cases are shown by the open circles.) Likewise, the participation ratio obtained numerically from the distribution of the positions visited during the localized periods only is quite close to the theoretical prediction (14), provided that $q$ is not too small for the same reason as above (Fig. 2d).
@@
Intermittent localization can be characterized by the lifetime of the localized periods, denoted as $t_{loc}$ and which is illustrated in Figs. 2a-b. This time is a random variable whose distribution $p(t_{loc})$ is computed numerically in Fig. 3a for several values of $q$. All the distributions are very well fitted by exponentials,
@@
practically over the whole range of lifetimes. Despite the non-Markov nature of the dynamics, $p(t_{loc})$ is remarkably simple and similar to the statistics of a Markov process, such as the escape from a potential well. In contrast, the durations $t_{deloc}$ of the delocalized periods follow a very different distribution. In Fig. 3b, it is well fitted at large times by an inverse power-law
$\rho(t_{deloc})\sim t_{deloc}^{-\alpha}$ (19)
@@
It follows from the analysis of Figs. 3a-b that intermittent localization is characterized by a finite mean localization time $\langle t_{loc}\rangle=\tau_{0}$, whereas the mean de-localization time $\langle t_{deloc}\rangle=\int_{0}^{\infty}dt\ t\rho(t)$ is infinite. The latter property is a consequence of the inequality $\alpha<2$. Hence very long trajectories will be dominated by very long diffusive segments of diverging durations interspersed by short localized periods, making the overall time-averaged occupation probability of the origin vanish asymptotically (when one does not distinguish between the two phases of motion). In this sense, the localized phase is “metastable”.
@@
The mean lifetime $\tau_{0}$ of the localized periods is determined by fitting the numerical distributions of Fig. 3a with Eq. (18). Figure 4a shows the variations of $\tau_{0}$ vs. $q$ at fixed $\gamma=0.7$, for several choices of $\Delta$. Interestingly, $\tau_{0}$ sharply increases with $q$ and rapidly becomes much larger than $\Delta$. The simulations suggest two possible scenarios: intermittent localization turns into full localization as $q$ crosses a critical value $q_{c}^{(I)}$ where the mean localization time diverges; or, the mean localization time is always finite but becomes extremely large for $q$ larger than a crossover value $q_{c}^{(I)}$. We shall come back to these scenarios in Section VII. Figure 3b displays the ’phase diagram’ for $\delta=0.7$: above $\Delta_{c}^{(I)}(q)$, the walker is localized in practice; below this value, it is intermittently localized or diffusive.
@@ IV.3 Power-law decaying memory with $1<\beta<2$
Let us retake the kernel (15), but in the range $1<\beta<2$. The mean forgetting time $\langle\tau\rangle$ in Eq. (16) is still infinite, while the effective number of remembered visits $C(t)$ in Eq. (3) saturates to a finite value at large $t$. Based on the numerical results obtained by varying $\beta$ within this interval, we speculate that this case is qualitatively similar to the exponential kernel of the previous section. We were not able to observe fully localized solutions in this range, contrary to the case $0\leq\beta\leq 1$. As exemplified by the numerical results in Figs. 5a-c for $\beta=1.6$, the system exhibits intermittent localization again with an occupation probability $P_{0}$ for the localized phase quite close to perfect case given Eq. (9). Meanwhile, the duration of the localization periods is still distributed exponentially and the mean localization time $\tau_{0}$ sharply increases with $q$.
@@ V Linear stability analysis of the localized solution
In this Section, we attempt to explain some of these results by studying Eq. (II) beyond the steady state regime. We assume that the positions $X_{t^{\prime}}$ and $X_{t}$ are still weakly correlated and use the factorization in Eq. (5)-(6), but keeping the time dependence of the occupation probabilities. The master equation (II) becomes
@@
Let us look for solutions taking the form of separation of variables, $g_{n}(t)=h(t)f_{n}$, and substitute this ansatz into Eq. (V). Dividing by $h(t)f_{n}$ on each side, separating the time-dependent part from the space-dependent one and setting them equal to a constant $\lambda$ independent of $n$ and $t$, one obtains
@@
where the amplitudes $a_{\lambda}$ depend on the initial condition, which is such that $P_{n}^{(st)}+g_{n}(t=t_{0})\geq 0$ for all $n$. In most of the following, we drop the indices $(\lambda)$ for brevity in the notation, and $h(t=0)=1$ by convention.
@@ V.1.1 Antisymmetric solutions
A first kind of solutions are anti-symmetric and denoted as $f_{n}^{(a)}$, characterized by $f^{(a)}_{0}=0$. In this case Eq. (24) is homogeneous for all $n$ and reads
@@ V.1.2 Symmetric solutions
The second type of solutions are symmetric, $f_{-n}=f_{n}$, and denoted as $f_{n}^{(s)}$. Eq. (24) must be considered in full with a non-zero $f_{0}^{(s)}$ that can be set to 11 (or some normalizing factor) without restricting generality. In this case, Eq. (26) becomes
@@
The poles of an arbitrary function $\widetilde{f}^{(s)}(\epsilon)$ in the complex plane, at $\epsilon=\{\epsilon_{1},\epsilon_{2},$...\}, correspond to exponential tails $f_{n}^{(s)}\sim c_{1}\epsilon_{1}^{-n}+c_{2}\epsilon_{2}^{-n}+$... for $n>0$ large enough, where the coefficient $c_{1},c_{2},$... depend on $(q,\gamma,\lambda^{\prime})$. Assume that there is a pole, e.g., $\epsilon_{1}$, with $|\epsilon_{1}|<1$. To ensure that $f_{n}^{(s)}$ does not diverge at large $n$, this pole must be canceled by imposing $c_{1}=0$. Another type of divergence can be caused by the presence of a double pole $\epsilon_{1}=\epsilon_{2}=1$ (implying that $f_{n}^{(s)}$ grows as $n$ at large $n$). In this case, one needs to impose $c_{1}=0$ as well. Solving $c_{1}(\lambda^{\prime})=0$ gives the sought eigenvalue $\lambda^{\prime}$ as function of $q$ and $\gamma$, and thus $\lambda$ via Eq. (29).
@@
$\widetilde{P}^{(st)}(\epsilon)=\sum_{n=1}^{\infty}\epsilon^{n}P_{n}^{(st)}=\frac{(1-\gamma)\epsilon P_{0}^{(st)}}{a-\epsilon}\,.$ (39)
Recall that $(r_{1},r_{2})$ depend on $\lambda^{\prime}$ and are given by Eq. (32). There are two cases:
$\bullet$ Case $|\lambda^{\prime}|<1$. In this case we can set,
@@
with $0<k<\pi$. This means that $r_{1},r_{2}$ are complex conjugates with $|r_{1}|=|r_{2}|=1$ like in the anti-symmetric solution. None of the 3 poles $\{r_{1},r_{2},a\}$ in Eq. (37)-(39) has a modulus smaller than 1 [$a>1$ from Eq. (11))], and no double pole exits. Therefore no pole needs to be canceled. For $n>0$ and large enough, the eigenfunctions are of the form
@@
$\bullet$ Case $\lambda^{\prime}\geq 1$. In this case $r_{1,2}$ are reals, with $r_{1}<1$ and $r_{2}>1$, or $r_{1}=r_{2}=1$ (double pole). As explained above, one must cancel the pole $r_{1}$ in Eq. (37), which is done by imposing $R(\epsilon=r_{1},\lambda^{\prime})=0$. The eigenvalue is thus root of the equation
@@
for all $0<q<1$ and $0<\gamma<1$. With the help of Mathematical, one can check that this solution to Eq. (42) is unique, independently of $(q,\gamma)$. The solution $\lambda^{\prime}=1$ implies that $r_{1}=r_{2}=1$ from Eq. (32), therefore Eq. (37) has a double pole and Eq. (42) guarantees that it is canceled. In this marginal case, corresponding to $k\to 0$ in Eq. (40), the eigenstate is a bound state since only the exponentially decaying mode remains, or

### v3

@@ III.2 Localization with $\gamma\neq 0$
Inserting these expressions into Eq. (II) gives the nonlinear equation,
@@
This equation can be exactly solved for the Fourier transform of $P_{n}^{(st)}$, where the occupation probability of the origin $P_{0}^{(st)}$ is considered as a constant that is finally obtained self-consistently. The analytical expression for the occupation probability at the origin was obtained in [20] and reads
@@
A quick analysis of Eq. (12) shows that $P_{0}^{(st)}>0$ for all $q>0$ and $\gamma>0$. Consequently, the localization length $\xi=(\ln a)^{-1}$ is always finite and the walker localizes for any non-zero parameter values $q$ and $\gamma$, a fact that was checked in numerical simulations of the model [20]. In addition, the agreement of Eqs. (12)-(13) with numerical simulations was very good (as also illustrated by Fig. 1b below), suggesting that the de-correlation approximation above might be exact. With the solution (12)-(13) at hand, it is also possible to calculate the participation ratio, a quantity of interest to characterize localization in classical [42] and quantum systems [43, 44, 45]. Here, it is defined as the probability that two independent walkers occupy the same site,
$PR=\sum_{n=-\infty}^{\infty}P_{n}^{2}\,$, (16)
and it is given by the expression
@@
Hence $PR$ is finite and does not decay to zero with the system size as in ordinary diffusion.
Localization in this context can be interpreted as a phenomenon of spatial learning, where the walker “learns” to exploit the resource patch at the origin instead of diffusing far away from it. As shown below in Fig. 1b, localization is stronger ($P_{0}^{(st)}$ larger) if the resource patch is more attractive ($\gamma$ large) or memory is more frequently used ($q$ large). Interestingly, on $3d$ lattices, the same de-correlation approximation predicts that a localization/delocalization phase transition takes place at $q=q_{c}$ with $q_{c}$ a non-zero critical value, in agreement with simulations [21]. In addition, the localization length $\xi$ diverges near $q_{c}$ with the same critical exponents as in the self-consistent theory of Anderson localization [20, 21]. If the n.n. random walk is replaced by a Lévy flight with sufficiently small index, a localization/delocalization transition takes place on the $1d$ lattice as well [20]. In the $1d$ problem studied here, $q_{c}=0$ and there is no de-localized phase.
@@
We now address the case of a general memory kernel $\pi_{t,t^{\prime}}$ and consider a resource site of attractiveness $\gamma>0$. If we apply the same approximations (8)-(III.2) to Eq. (II), we obtain Eq. (11) again. This is simply due to the fact that the normalization condition $\sum_{t^{\prime}=0}^{t}\pi_{t,t^{\prime}}=1$ is valid for all kernels. Hence, under the same assumptions, the results (12)-(17) hold for any kernel, not just the perfect memory case. This, of course, cannot be correct if memory is very short-range, with $\pi_{t,t^{\prime}}$ very peaked near the present time $t$, i.e., $F(\tau)$ tends rapidly to 0 in Eq. (2). In that case, we expect the walker to behave roughly as a standard random walk with a modified diffusion coefficient, as shown in [34] in the absence of impurity. No stationary localized solutions should be expected in principle.
To build intuition on the behavior of the system in the different cases, we have run Monte Carlo simulations generating stochastic trajectories with the rules of the model. Appendix A gives details on the generation of the random times $t^{\prime}$. We have considered the following choices for the function $F(\tau=t-t^{\prime})$.
Figure 1: (a) A localized trajectory $X_{t}$ vs. t with power-law memory decay, for $\beta=1$, obtained from a numerical simulation with $q=0.1$ and $\gamma=0.7$. (b) Occupation probability of the origin, $P_{0}$, as a function of $q$ for various values of $\gamma$. Solid (blue) lines are given by Eq. (12) and symbols by Monte Carlo simulations ($t=10^{6}$) corresponding to different memory decay exponents $\beta\in[0,1]$.
@@ IV.1 Slowly decaying memory
In the first class of examples, the memory slowly decays as a power-law, and not faster than $1/\tau$. This type of law was observed in experiments on human subjects (the Ebbinghaus’ forgetting curve [46]) or inferred in the wild from the trajectories of large herbivores [32]. We denote,
@@
where $\tau=0,1,$... In this range of exponent values, the mean forgetting time, defined as [34]
$\langle\tau\rangle=\lim_{t\to\infty}\frac{\sum_{t^{\prime}=0}^{t}(t-t^{\prime})F(t-t^{\prime})}{\sum_{t^{\prime}=0}^{t}F(t-t^{\prime})}\,$, (19)
is infinite. In addition, the effective number of visits that are remembered by the walker up to time $t$ [Eq. (3)] keeps increasing unbounded, albeit more slowly than $t$. Hence, memory is quite long-range for these kernels and the perfect memory (non-decaying) is recovered with $\beta=0$.
We display in Fig. 1a a typical trajectory $X_{t}$ vs. $t$ with $\beta=1$, showing a clear stationary localized behavior at late times, where the particle is never very far away from the origin. Figure 1b shows the occupation probability $P_{0}$ computed at $t=10^{6}$ as a function of $q$ for various values of $\gamma$. In each case, the curves corresponding to different values of $\beta$ in the interval $[0,1]$ practically fall on top of each other and are in very good agreement with the theoretical prediction (12) of the de-correlation approximation. Hence, the applicability range of this approximation is not limited to the perfect memory studied in [20] and extends to kernels with sufficiently slow forgetting as well.
@@ IV.2 Exponentially decaying memory
Two typical trajectories are displayed in Figs. 2a-b and a new type of behavior can be observed: the walker is not completely localized nor diffusive but stochastically alternates between these two types of motion. Fig. 2b, for instance, shows a very long diffusive phase at late times. We have identified the two phases using a simple algorithm described in Appendix B and have analyzed them separately. Notably, the occupation probability $P_{0}$ at the origin computed from averaging only over the localized parts of many trajectories, shown in Fig. 2c, is practically undistinguishable from $P_{0}^{(st)}$ in Eq. (12). In comparison, the occupation probability $P_{0}$ in the de-localized phase is very low. (Note that if $q$ is too small, the identification of the two phases becomes difficult and cannot be done in a reliable way; these cases are labeled with open circles.) Likewise, the inverse participation ratio obtained numerically from the distribution of the positions visited during the localized periods only is quite close to the theoretical prediction (17), provided that $q$ is not too small for the same reason as above (Fig. 2d).
@@
Intermittent localization can be further characterized by the lifetime of the localized periods, denoted as $t_{loc}$ and which is illustrated in Figs. 2a-b. This time is a random variable whose distribution $p(t_{loc})$ is computed numerically in Fig. 3a for several values of $q$. All the distributions are very well fitted by exponentials,
@@
practically over the whole range of lifetimes $t_{loc}$. Despite the non-Markov nature of the dynamics, $p(t_{loc})$ is remarkably simple and similar to the statistics of a Markov process, such as the escape from a potential well. In contrast, the durations $t_{deloc}$ of the delocalized periods follow a very different distribution. In Fig. 3b, it is well fitted at large times by an inverse power-law
$\rho(t_{deloc})\sim t_{deloc}^{-\alpha}\,$, (22)
@@
It follows from the analysis of Figs. 3a-b that intermittent localization is characterized by a finite mean localization time $\langle t_{loc}\rangle=\tau_{0}$, whereas the mean de-localization time $\langle t_{deloc}\rangle=\int_{0}^{\infty}dt\ t\rho(t)$ diverges. The latter property is a consequence of the inequality $\alpha<2$. Hence very long trajectories will be dominated by very long diffusive segments of diverging durations interspersed by relatively short localized periods, making the overall time-averaged occupation probability of the origin vanish asymptotically (when one does not distinguish between the two phases of motion). In this sense, the localized phase is “metastable”.
@@
The mean lifetime $\tau_{0}$ of the localized periods is determined by fitting the numerical distributions of Fig. 3a with Eq. (21). Figure 4a shows the variations of $\tau_{0}$ vs. $q$ at fixed $\gamma=0.7$, for several choices of $\Delta$. Interestingly, $\tau_{0}$ sharply increases with $q$ and becomes larger than $\Delta$ by orders of magnitude. The simulations suggest two possible scenarios: intermittent localization turns into full localization as $q$ crosses a critical value $q_{c}^{(I)}$ where the mean localization time diverges; or, the mean localization time is always finite but becomes extremely large for $q$ larger than a crossover value $q_{c}^{(I)}$. We shall come back to these scenarios in Section VII. Figure 3b displays the ’phase diagram’ showing the crossover region with dots for $\delta=0.7$: above $\Delta_{c}^{(I)}(q)$, the walker is localized in practice; below this value, it is intermittently localized or diffusive.
@@ IV.3 Power-law decaying memory with $1<\beta<2$
Another situation of relatively fast forgetting corresponds to the kernel (18) in the range $1<\beta<2$. The mean forgetting time $\langle\tau\rangle$ in Eq. (19) is still infinite in that case, but the effective number of remembered visits $C(t)$ in Eq. (3) saturates to a finite value at large $t$. Based on the numerical results obtained by varying $\beta$ within this interval, we speculate that this case is qualitatively similar to the exponential kernel of the previous section. We were not able to observe fully localized solutions in this range, contrary to the case $0\leq\beta\leq 1$. As exemplified by the numerical results in Figs. 5a-b-c for $\beta=1.6$, the system exhibits intermittent localization again with an occupation probability $P_{0}$ for the localized phase quite close to the perfect case given Eq. (12). Meanwhile, the duration of the localization periods is still distributed exponentially and the mean localization time $\tau_{0}$ sharply increases with $q$.
@@ V Linear stability analysis of the localized solution
In this Section, we attempt to explain some of these results by studying Eq. (II) beyond the steady state regime. We assume that the positions $X_{t^{\prime}}$ and $X_{t}$ are still weakly correlated and use the factorization in Eq. (8)-(9), but keeping the time dependence in the occupation probabilities. The master equation (II) becomes
@@
Let us look for solutions taking the form of separation of variables, $g_{n}(t)=h(t)f_{n}$, and substitute this ansatz into Eq. (V). Dividing by $h(t)f_{n}$ on each side, separating the time-dependent part from the space-dependent one and setting them equal to a constant $\lambda$ independent of $n$ and $t$, we obtain
@@
where the amplitudes $a_{\lambda}$ depend on the initial condition, which is such that $P_{n}^{(st)}+g_{n}(t=t_{0})\geq 0$ for all $n$. In most of the following, we drop the indices $(\lambda)$ for brevity in the notation, and we set $h(t=0)=1$ by convention.
@@ V.1.1 Antisymmetric solutions
A first kind of solutions, denoted as $f_{n}^{(a)}$, are anti-symmetric and characterized by $f^{(a)}_{0}=0$. In this case, Eq. (27) is homogeneous for all $n$ and reads
@@ V.1.2 Symmetric solutions
The second type of solutions are symmetric, $f_{-n}=f_{n}$, and denoted as $f_{n}^{(s)}$. Eq. (27) must be considered in full with a non-zero $f_{0}^{(s)}$ that can be set to $1$ (or some normalizing factor) without restricting generality. In this case, Eq. (29) becomes
@@
The poles of an arbitrary function $\widetilde{f}^{(s)}(\epsilon)$ in the complex plane, at $\epsilon=\{\epsilon_{1},\epsilon_{2},$...\}, correspond to exponential tails $f_{n}^{(s)}\sim c_{1}\epsilon_{1}^{-n}+c_{2}\epsilon_{2}^{-n}+$... for $n>0$ large enough, where the coefficients $c_{1},c_{2},$... depend on $(q,\gamma,\lambda^{\prime})$. Assume that there is a pole, e.g., $\epsilon_{1}$, with $|\epsilon_{1}|<1$. To ensure that $f_{n}^{(s)}$ does not diverge at large $n$, this pole must be canceled by imposing $c_{1}=0$. Another type of divergence can be caused by the presence of a double pole $\epsilon_{1}=\epsilon_{2}=1$ (implying that $f_{n}^{(s)}$ grows as $n$ at large $n$). In this case, one needs to impose $c_{1}=0$ as well. Therefore, solving $c_{1}(\lambda^{\prime})=0$ gives the sought eigenvalue $\lambda^{\prime}$ as a function of $q$ and $\gamma$, and finally $\lambda$ via Eq. (32).
@@
$\widetilde{P}^{(st)}(\epsilon)=\sum_{n=1}^{\infty}\epsilon^{n}P_{n}^{(st)}=\frac{(1-\gamma)\epsilon P_{0}^{(st)}}{a-\epsilon}\,$, (42)
from the known solution (13). Recall that $(r_{1},r_{2})$ depend on $\lambda^{\prime}$ and are given by Eq. (35). There are two cases:
$\bullet$ Case $|\lambda^{\prime}|<1$. In this case, we can set
@@
with $0<k<\pi$. This means that $r_{1},r_{2}$ are complex conjugates with $|r_{1}|=|r_{2}|=1$ like in the anti-symmetric solution. None of the 3 poles $\{r_{1},r_{2},a\}$ in Eq. (40)-(42) has a modulus smaller than 1 [$a>1$ from Eq. (14)], and a double pole does not exist. Therefore, no pole needs to be canceled. For $n>0$ and large enough, the eigenfunctions are of the form
@@
$\bullet$ Case $\lambda^{\prime}\geq 1$. In this case $r_{1,2}$ are reals, with $r_{1}<1$ and $r_{2}>1$, or $r_{1}=r_{2}=1$ (double pole). As explained above, one must cancel the pole $r_{1}$ in Eq. (40), which is done by imposing $R(\epsilon=r_{1},\lambda^{\prime})=0$. The eigenvalue is thus a root of the equation
@@
for all $0<q<1$ and $0<\gamma<1$. With the help of the software Mathematica®, one can check graphically with many pairs of parameters $(q,\gamma)$ that Eq. (45) does not admit other real solutions. Hence, the solution $\lambda^{\prime}=1$ is unique, independent of $(q,\gamma)$, and implies that $r_{1}=r_{2}=1$ from Eq. (35). In other words, Eq. (40) has a double pole and Eq. (45) guarantees that it is canceled. In this marginal case, corresponding to $k\to 0$ in Eq. (43), the eigenstate is a bound state since only the exponentially decaying mode remains, or